OpenStudy (anonymous):

One of the friends may help me to translate this! Its Arabic :) مجذور اذا زد نا علیه عشره کان للمجتمع جذر او نقصانها منه کان للباقی جذر

OpenStudy (anonymous):

I think it means to find a number \(x\) such that\[\sqrt{x^2+10}=a\]\[\sqrt{x^2-10}=b\]

ganeshie8 (ganeshie8):

google translate thinks it is persian : `Inasmuch square root Llmjtm he was ill against Shrh Kahn Kahn my deficiencies square Llbaqy`

OpenStudy (anonymous):

see what it gave me from Arabic to English: If we increase the roots of it it was ten in the community root or incompleteness of it was for the rest of the root I have no Idea :)))))

ganeshie8 (ganeshie8):

@ikram002p

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

No, its Arabic, I know because I speak Persian and I've studies Arabic.

OpenStudy (anonymous):

*studied

OpenStudy (anonymous):

well, LYRA_Z where are you from? I'm from Iran by the way

OpenStudy (anonymous):

Me too!

OpenStudy (anonymous):

So what is your question exactly?

OpenStudy (anonymous):

(I don't live in Iran though bth)

OpenStudy (anonymous):

well this is interesting, This is one of the seven problems of Sheykh Bahayee, from his book "Kholasat-ol-Hesab" and I don't what exactly wants in this question

OpenStudy (anonymous):

don't know*

OpenStudy (anonymous):

He has said in the introduction of problems that these problems are unsolved until now, but now everybody with some high school information about math can solve them :)

OpenStudy (anonymous):

I can upload the questions for you, If you are interested :)

OpenStudy (anonymous):

Yes, that would be nice!

OpenStudy (md152727):

I can speak arabic but can't solve this problem , sorry

OpenStudy (anonymous):

I don't want you to solve it, just translate the problem :)

OpenStudy (anonymous):

Interesting...I'll try to solve them on my free time! Let me see what I can do to translate it...although it seems an Arabic speaker has come to help!

OpenStudy (er.mohd.amir):

@mukushla Good Q in pdf i try to first understand it then try to solve.

OpenStudy (md152727):

I am native- arabic speaker but can't help in math I am still in 9th grade, I am 11

OpenStudy (anonymous):

@MD152727 But can you translate it?

OpenStudy (md152727):

Yes

OpenStudy (anonymous):

Well that's what @mukushla wants, just a translation!

OpenStudy (md152727):

Oh Ok! Easy!

OpenStudy (anonymous):

Woah wait a minute the pdf you sent has Farsi!

OpenStudy (md152727):

What is Farsi?

OpenStudy (anonymous):

@LYRA_Z I think the translation in the PDF is wrong

OpenStudy (anonymous):

Farsi is a Persian Language.

OpenStudy (anonymous):

@mukushla oh....are you sure?

OpenStudy (md152727):

ok, Never heard of it

OpenStudy (anonymous):

yeah, I'm sure, because the question is trivial in that way

OpenStudy (md152727):

مجذور اذا زد نا علیه عشره کان للمجتمع جذر او نقصانها منه کان للباقی جذر : In English: A root if added by 10 will equal root or subtracted from it which will equal roots

OpenStudy (anonymous):

oh...

OpenStudy (md152727):

Man! I can't understand this math

OpenStudy (anonymous):

Thanks anyway @MD152727 :)

OpenStudy (md152727):

Your welcome

OpenStudy (anonymous):

Is that all?

OpenStudy (anonymous):

No, I'm still stuck :)))

OpenStudy (anonymous):

On the problem or the translation or both?

OpenStudy (anonymous):

Now somebody put it in mathematical words XD

OpenStudy (anonymous):

Not sure how to help with that...XD What exactly do you mean by "mathematical words"?

OpenStudy (anonymous):

well set up equations for the problem I mean :)

OpenStudy (anonymous):

setting up the*

OpenStudy (anonymous):

hmmm....don't think I can be much help...

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Good luck!

OpenStudy (ikram002p):

its weird as "مجذور" is Persian, syntax in Arabic mean no since ... where did u got it from?

OpenStudy (anonymous):

I have uploaded a PDF ikram, Its in there :)

OpenStudy (ikram002p):

oh!

OpenStudy (ikram002p):

are you sure the origin writes of the problems i Arabic ? seems like broken Arabic *It's annoying* :O

OpenStudy (ikram002p):

ok... i'll just translated word by word mixing Persian with Arabic lol xD

OpenStudy (anonymous):

I don't know exactly, did you look at PDF? thanks ikram

OpenStudy (ikram002p):

it says a square if we added 10 to it then the whole root and it's sum to 10 can be a root to an integer (this means as u said solve for integer \(\sqrt{x^2+10}\) ) but if it was incomplete square ( mathematically means \(x^2-10\)) then it would be root for an integer.(same : solve for integer \(\sqrt{x^2-10}\) )

OpenStudy (ikram002p):

can you read Persian/ Arabic ? so i would note were you would misunderstood the question ?

OpenStudy (anonymous):

do you mean translate what has written in persian in the PDF?

OpenStudy (ikram002p):

مجذور = can hold both root and square meanings (In its syntax) but been used as Square in mathematics in ancient time . للمجتمع=It means community so it sound like the writer in not Arabian native as its build wrong if he/she mean "summation or whole square " ... the rest would make sense...

OpenStudy (anonymous):

thanks for your complete answer ikram :) I got it :)

OpenStudy (ikram002p):

np =)

OpenStudy (anonymous):

Two of the seven problems are interesting actually, rest are not good ;) One is the following system\[x^2+x+2=a^2 \\ x^2-x-2=b^2\]for real \(x\) and integers \(a\) and \(b\) and the second the system\[x^2+y^2+z^2=a^2 \\ \frac{x^2}{y^2}=\frac{y^2}{z^2}=\frac{z^2}{x^2}\]

OpenStudy (ikram002p):

quadratic diophantine equations.. i would like to se how its solved by that ancient method =)

OpenStudy (anonymous):

:)

OpenStudy (mathmale):

Please go to translate.google.com and use that tool to translate this Arabic.

Nnesha (nnesha):

yes?

Nnesha (nnesha):

No hablo arabic :(

OpenStudy (anonymous):

@mathmale Sorry but actually this question should be in mathematics, because my main problem was to put it in mathematical phrases!

OpenStudy (anonymous):

Anyway, thanks for closing the question, It was solved!

Nnesha (nnesha):

2nd and 3rd comment on this post) he already mentioned about google translate... he used that and didn't understand the question i guess. and google translate doesn't really give accurate translation all the time.